OpenStudy (anonymous):

how do i solve this log2 3+log2 x=3

OpenStudy (arbershabani97):

then log2(3x)=3 2^3=3x 8=3x x=8/3

OpenStudy (anonymous):

ok great :)thats what i got

OpenStudy (anonymous):

i need help with another one

OpenStudy (campbell_st):

well using log laws log(ab) = log(a) + log(b) \[\log_{2}(3 \times x) = 3\] which when simplified is \[\log_{2}(3x) = 3\] so now raise each side as a power of the base \[2^{\log_{2}(3x)} = 2^3\] which simplifies to 3x = 8 now just solve for x

OpenStudy (arbershabani97):

ok just write it

OpenStudy (anonymous):

3^2x+2=8 solve for x..

OpenStudy (anonymous):

i got -0.15

OpenStudy (arbershabani97):

3^2x=6 ln3^2x=ln6 2x ln3=ln6 2x=ln6/ln3 2x=1.630929754 x=0.815464876

OpenStudy (anonymous):

the exponent is 3^(2x+2)

OpenStudy (arbershabani97):

oh, i thought the +2 was out of exponent wait a second

OpenStudy (anonymous):

thanks :)

OpenStudy (arbershabani97):

2x+2=ln8/ln3 2x+2=1.892789261 x=-0.053605369

OpenStudy (anonymous):

yay perfect :D

OpenStudy (anonymous):

thank you :)

OpenStudy (arbershabani97):

you're welcome :)